OpenStudy (anonymous):

How do I solve this (if the area does cross the y-axis)? I need to find the area of the shaded area (area between the following two functions): upper boundary: y= 4-3x lower boundary: y = 3x^2 - 3x +1 Thank you!

OpenStudy (anonymous):

First set 4-3x = 3x^2 -3x +1 to get your intercepts. Once you know the x values you can sub them into your definite integral.

OpenStudy (anonymous):

Your second step will look like this: \[\int\limits_{-1}^{1} (4-3x) - (3x^2-3x+1) dx\]

OpenStudy (anonymous):

Good luck!

OpenStudy (anonymous):

Yay! I get (2) - (-2) = 4 ae. Thank you!!